# Extracting the singularity of the logarithmic partition function

### v1

@@ Abstract
In principle, the logarithm of the partition function and its derivatives can be used to rigorously determine the nature of phase transitions. However, in practice, this approach is often ineffective in numerical calculations, since resolving the singularities of the logarithmic partition function requires extremely high-precision data—especially when distinguishing between weakly first-order phase transitions and deconfined quantum critical points. Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the leading volume-law term. This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE). We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems. Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in $J\text{-}Q$ models, which were initially thought to be continuous deconfined quantum phase transitions. The discontinuity of these transitions is challenging to detect using traditional local physical quantities. This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
@@
Introduction.— Quantum phase transitions, spontaneous symmetry breaking, and the universality classes of critical phenomena are central concepts in condensed matter and statistical physics [1, 2, 3]. The type of phase transition is well defined by the singularity of free energy and its derivative. For a first-order phase transition, the first-order derivative of free energy is discontinuous. Otherwise, if its higher-order derivative is discontinuous, it is named a continuous phase transition. In principle, if the partition function can be calculated exactly, the phase transition will be fixed strictly according to the singularity. However, in practical numerical simulations, directly detecting such singularities through the free energy and its derivatives is difficult, as it requires extremely high-precision data. This challenge is particularly pronounced when distinguishing between weakly first-order phase transitions and deconfined quantum critical points  [4, 5, 6, 7, 8, 9, 10]. A major reason for this difficulty is that the singular contribution is typically overshadowed by a much larger, non-singular term that scales with the system volume. How to extract the singularity remains a significant challenge in theoretically computational physics.
@@
Considering a quantum phase transition ($\beta\rightarrow\infty$), the singularity of free energy can be replaced by focusing on the logarithmic partition function. According to the finite-size scaling hypothesis [11, 12], an observable which is singular at $g_{c}$ (the parameter $g$ of Hamiltonian at the phase transition point) scales with the system size close to $g_{c}$ as a power of $L$ multiplied by a non-singular function of the ratio $\xi/L$, where $\xi$ is the correlation length and $L$ is the system size. Any singular quantity (not necessarily divergent) should thus be of the form:
$Q(\Delta g,L)=L^{\sigma}f(\Delta gL^{1/\nu})$ (1)
@@
where $\Delta g=g-g_{c}$, $\nu$ is the critical exponent of the correlation length and $\sigma$ is an undetermined power. Since the singular part of the free energy is scale-invariant ($\sigma=0$[13, 11], the logarithmic partition function can be written in the following form:
$\ln Z(g)=a(g)L^{d}+f(\Delta gL^{1/\nu})$ (2)
@@
where $Z$ is a partition function, $d$ is the spatial dimension and $a(g)L^{d}$ represents the leading term obeying a volume law. Here, we see that this dominant volume law term often masks the singular contribution of $f$, which explains the difficulty in numerically detecting the phase transition through such singularities.
@@
An approach to extract the singular term $f$ is to construct a quantity in which the volume-law contributions cancel out, leaving only the singular part. Inspired by the method used to calculate topological entanglement entropy [14, 15], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\rightarrow\infty$) isolates the singular term $f$, with the dominant volume-law contributions canceled out. Specifically, we argue that Eq. (2) can be rewritten as $\ln Z(2\beta)=2a(g)L^{d}+f$ and $2\ln Z(\beta)=2a(g)L^{d}+2f$. Coincidentally, we find that this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [16, 17, 18, 19, 20, 21, 22]:
@@
This definition can be generalized to an arbitrary order $n$, leading to the $n$-th order RTE: $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$. For simplicity, we focus on the 2nd-order RTE ( i.e., $n=2$) throughout this paper.
@@
In this work, assisted by our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [23, 24, 25, 26, 27, 28, 29], we demonstrate the validity of our conjecture above: the RTE is indeed a dimensionless quantity that captures only the singular contribution. Furthermore, we observe a more pronounced singularity in the derivative of the RTE (DRTE) near the quantum critical point (QCP). We propose universal scaling relations for both the RTE and DRTE, which are confirmed by data collapse analyses performed near (2+1)D QCPs belonging to the O($N$) universality classes, with $1\leq N\leq 3$. More interestingly, since the leading non-singular term of $\ln Z$ is canceled out in the DRTE, it becomes a particularly powerful tool for detecting weakly first-order phase transitions even in normal sizes, which are notoriously difficult to resolve using conventional numerical approaches.
@@
Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$. Additionally, we explore first-order transitions, with a particular focus on the challenging weakly first-order phase transitions (e.g., some controversial deconfined quantum critical points). Detailed descriptions of the model Hamiltonians will be provided in the following sections.
@@
As shown by the definitions above, the 2nd-order Rényi thermal entropy can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. The partition functions can be calculated using the bipartite reweight-annealing algorithm [23, 24] in conjunction with stochastic series expansion (SSE) QMC simulations [30, 12, 31, 32, 33, 34, 35]. Furthermore, following a procedure similar to that used in calculating the derivative of Rényi entanglement entropy and negativity [23, 36], the derivative of the RTE with respect to a general parameter $J$ can be written as:
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
In this expression, the first average is evaluated on the ensemble corresponding to $Z(n\beta)$, while the second is based on $Z(\beta)$. In our simulations, the inverse temperature $\beta$ is set to be linearly proportional to the lattice size $L$ (i.e., $\beta=L$) in order to approach the quantum critical regime. Further methodological details are provided in the Supplementary Materials (SM).
@@
Finite-size scaling form.— Considering a zero-temperature system and assuming the ground-state degeneracy is $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$+$f(\Delta gL^{1/\nu})$, where $f$ represents the critical corrections. As shown in Fig.1, $S^{(2)}$ can be used to probe both the degeneracy of discrete symmetry-breaking and the tower of excited states (TOS) associated with continuous symmetry-breaking.
@@
More interestingly, as shown in Fig. 1, we observe that the curves for different system sizes cross at the critical point, which implies that $S^{(2)}$ is indeed a dimensionless quantity, consistent with our argument in the introduction. Near the critical point, the logarithmic partition function can be decomposed into a non-singular part $a(J)L^{d}$ and a dimensionless singular term $f(\Delta gL^{1/\nu})$. The non-singular contribution is the leading term, which makes it difficult to extract critical information directly from the partition function. As shown in Fig. 1, the crossings of curves for different sizes indicate that the RTE is independent of the system size. This provides strong numerical support for our argument that the leading volume-law contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel out in the definition of RTE, leaving only the dimensionless singular part. Additionally, we observe that the DRTE exhibits a stronger divergence near the QCP.
@@
According to the finite-size scaling hypothesis  [11, 12], any singular quantity near a critical point should obey a scaling form as given in Eq.(1). In particular, the dimensionless RTE ($\sigma=0$), is expected to follow the scaling form:
$S^{(n)}(\Delta g,L)=f(\Delta gL^{1/\nu}).$ (5)
@@
$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=f^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
where $f^{\prime}(x)$ is $\partial f(x)/\partial x$ and $x$ is a general parameter.
Figure 1: (a) The 2nd-order RTE and (c) its derivative, and (b) and (d) the data collapse analysis of the RTE and DRTE near the QCP of the bilayer Ising-Heisenberg model (with (2+1)D Ising universality class). We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Data of the dimerized anisotropic Heisenberg model with $\Delta=0.9$ (with (2+1)D O(2) universality class). We set $J_{2}=2.1035$, and take $J_{1}$ as the tuning parameter. (i–m) Data of the dimerized isotropic Heisenberg model with $\Delta=1$ (with (2+1)D O(3) universality class). We set $J_{2}=1$, and take $J_{1}$ as the tuning parameter. To achieve good data collapse, a subleading correction is also included: $Q(g,L)=L^{k}[f((g-g_{c})L^{1/\nu})+bL^{-1}]$.
In the following section, by expanding the scaling function $f(\Delta gL^{1/\nu})$ (or $f^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)d O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
Ising transition.— We consider the bilayer Ising-Heisenberg model with (2+1) D Ising transition  [37]. The Hamiltonian reads:
@@
in which $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ is the layer index. $-J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction. For $J/J_{z}\ll 1$, the ground state exhibits Ising order, breaking the $\mathbb{Z}_{2}$ symmetry. For $J/J_{z}\gg 1$, the ground state is in a spin-singlet states. The QCP separating two phases is at $J/J_{z}=3.045(2)$ and belongs to the (2+1)D Ising universality class [37].
We set $J=3.045$ and use $J_{z}$ as the tuning parameter. We then evaluate the 2nd-order RTE and its derivative in the vinicity of the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Fig. 1 (a) and (c). The RTEs of different sizes intersect at QCP, and their derivatives have a peak at QCP. We fit the scaling forms given by Eq. (5) and (6) to the data of RTE and its derivative separately. For the RTE data, we find $J_{z,c}=0.9999(1)$ and $\nu=0.62(1)$, while for the derivative of the DRTE data, we find $J_{z,c}=1.000(3)$ and $\nu=0.63(4)$. These results are consistent with the QCP $J_{z,c}=1$  [37] and the critical exponent $\nu=0.63012$ in the (2+1)D Ising universality class [38, 39, 37]. The universal scaling functions $f(\Delta gL^{1/\nu})$ and $f^{\prime}(\Delta gL^{1/\nu})$ obtained from the data collapse analysis are plotted in Fig. 1 (b) and (d).
@@
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction and $\Delta$ is anisotropy parameter. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $10<\Delta<$ 1, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $10<\Delta<$ 1/ $\Delta>1$ / $\Delta=1$  [40, 41, 42].
@@
The RTE and its derivative are evaluated around the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality). The results are shown in Fig. 1 (e,g) and (i,k), respectively. In each case, we fix $J_{2}$ and tune $J_{1}$ across the QCP, as described in the figure caption. Similar to the Ising transition case, we find that the RTE values for different lattice sizes intersect at the QCPs, and their derivatives exhibit peaks at the QCPs.
We fit the scaling forms given by Eq. (5) and (6) near the (2+1)D O(2) and O(3) QCPs separately. Around the O(2) QCP, we find $J_{2,c}=0.997(1)$ and $\nu=0.67(2)$ for the RTE data, and $J_{2,c}=0.999(1)$ and $\nu=0.66(2)$ for the DRTE data. These results are consistent with the known QCP and critical exponent: $J_{1,c}=1$ and $\nu=0.6703$ [38, 42].
Near the O(3) QCP, we find $J_{2,c}=0.523(1)$ and $\nu=0.70(1)$ for the RTE data, and $J_{2,c}=0.525(1)$ and $\nu=0.71(2)$ for the DRTE data. These results are also consistent with the known QCP and the critical exponent: $J_{1,c}=0.52337$ and $\nu=0.7073$ [38, 40, 41]. The universal scaling functions near O(2) and O(3) QCPs obtained from the data collapse analysis are shown in Fig. 1 (f), (h) and (g), (m) .
@@
Figure 2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$, and (b) and (c) the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model.
@@
First-order and weakly first-order transitions.— Having established the universal scaling form for (2+1)D QCPs, we then investigate first-order transitions, especially the weakly first-order phase transition presenting significant challenges for traditional local physical quantities.
@@
We first study the 2D checkerboard $J$-$Q$ model, which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [43]. The DRTE around this transition is plotted in Fig. 2 (a). A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side of the transition as $L\to\infty$. This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above). For comparison, we also calculate the derivative of logarithmic partition functions, $2\partial\ln Z(\beta)/\partial J$ and $\partial\ln Z(2\beta)/\partial J$, as shown in Fig. 2 (b) and (c). The curve seems much more continuous compared with DRTE, because the derivative of logarithmic partition function includes a much larger term obeying volume-law. The phase transition of this model is obviously first-order according to the previous works [43, 23], thus we see that the cuves in (b) and (c) have very small bends around the phase transition point, especially for larger sizes.
@@
The phase transition between AF Néel order and valence-bond solid in 2D quantum magnets can be (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [44, 45]. $J-Q_{2}$ and $J-Q_{3}$ models can realize the DQC in 2D quantum spin systems [46, 47]. This model has no sign problem and is, therefore, amenable to QMC simulations. Numerous QMC studies have shown that the phase transitions of the $J-Q$ model on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [46, 47, 48, 49, 50, 51]. However, the exponents of the $J-Q$ models do not satisfy the conformal bootstrap bounds of a $SO(5)$ critical point [52]. Recent studies have provided growing evidence suggesting that this DQC point is in fact a weakly first-order phase transition [53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. In the following, we demonstrate that DRTE can clearly and simply prove that the phase transition in the $J-Q_{2}$ and $J-Q_{3}$ systems is a first-order phase transition.

### v2

# Extracting the singularity of a logarithmic partition function to investigate quantum phase transitions

@@ Abstract
In principle, the logarithm of a partition function and its derivatives can be used to rigorously determine the nature of phase transitions. However, this approach is often ineffective in numerical calculations in practice, since resolving the singularity of a logarithmic partition function requires extremely high-precision data—especially when distinguishing weakly first-order phase transitions and deconfined quantum critical points. Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the analytic terms. This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE). We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems. Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in JQJ-Q models, which were initially thought to be continuous deconfined quantum phase transitions. The discontinuity of these transitions is challenging to detect using traditional local physical quantities. This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
@@
Introduction.— Quantum phase transitions, associated spontaneous symmetry breaking, and universality classes are central concepts in condensed matter and statistical physics [1, 2, 3]. The nature of a phase transition is determined by the singularity in the free energy or its derivatives: a first-order transition is marked by a discontinuity in the first derivative, while a continuous transition involves singular higher-order derivatives. In principle, an exact calculation of the partition function would allow one to identify the transition based on these singularities. However, in practice, detecting such features in numerical simulations is challenging, as it requires extremely high-precision data. This issue is particularly severe when distinguishing weakly first-order transitions from deconfined quantum critical points [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. A key difficulty arises because the singular contribution is often overwhelmed by a much larger, non-singular term that scales with the system volume. Extracting the singular part therefore remains a major challenge in computational physics.
@@
For a quantum phase transition (inverse temperature $\beta\rightarrow\infty$; in the realistic simulations, the finite-size system gap $\Delta_{L}\sim\frac{1}{L^{z}}$, $\beta\sim L^{z}$ is usually used for probing the ground state properties, in which $L$ is the system size and $z$ is the dynamical critical exponent [17] ), the singularity of the free energy can be equivalently studied through the logarithm of the partition function. According to the finite-size scaling hypothesis [18, 17], a singular observable $Q$ (not necessarily divergent) near the critical point $g_{c}$ scales with system size as:
$Q(\Delta g,L)=L^{\sigma}\tilde{S}(\Delta gL^{1/\nu})$ (1)
@@
where $\Delta g=g-g_{c}$, $\nu$ is is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function. Since the singular part of the free energy is scale-invariant ($\sigma=0$[19, 18], the logarithm of the partition function can be expressed as:
$\ln Z(g,\beta,L)=a(g)\beta L^{d}+\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}(\Delta gL^{1/\nu})$ (2)
@@
where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero or non-zero). The dominant volume-law term often obscures the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
@@
To isolate the singular term $\tilde{S}$, one can construct a quantity in which the dominant analytic contributions cancel out. Inspired by the method used to extract topological entanglement entropy [20, 21], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\sim L^{z}$) can serve this role. Specifically, we argue that Eq. (2) can be rewritten as $\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ and $2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$. Coincidentally, this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [22, 23, 24, 25, 26, 27, 28]:
@@
This definition generalizes naturally to arbitrary order $n$ : $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$. For clarity and simplicity, we focus on the second-order RTE ($n=2$) throughout this work.
@@
In this work, using our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [29, 30, 31, 32, 33, 34, 35], we validate the conjecture that the Rényi thermal entropy (RTE) is a dimensionless quantity capturing only the singular contribution. Moreover, its derivative (DRTE) exhibits a more pronounced singularity near the quantum critical point (QCP). We propose universal scaling forms for both RTE and DRTE, which are confirmed through data collapse analyses at (2+1)-dimensional (D) QCPs in the O($N$) universality classes, with $1\leq N\leq 3$. Notably, because the leading non-singular term in $\ln Z$ is canceled in DRTE, it provides a powerful diagnostic for weakly first-order phase transitions, even in relatively small systems (up to $L=56$), which are significant challenges to resolve using conventional numerical approaches.
@@
Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$. We also examine first-order transitions, with a focus on weakly first-order cases, including some debated deconfined quantum critical points. The model Hamiltonians are described in detail in the following sections.
@@
The second-order RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. We compute the partition functions using the bipartite reweight-annealing algorithm [29, 30] in combination with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 39, 40, 41]. Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [29, 42], the derivative of the RTE with respect to a general parameter $J$ is given by:
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively. While the RTE is calculated using the reweight-annealing QMC method, the derivative (DRTE) is directly obtained from the two expectation values in Eq. (4); see Supplementary Materials (SM) for details. Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
@@
Finite-size scaling form.—For a zero-temperature system with ground-state degeneracy $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$$\tilde{S}(\Delta gL^{1/\nu})$, where $\tilde{S}$ represents the critical corrections. As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the tower of excited states (TOS) associated with continuous symmetry-breaking.
@@
Near criticality, the logarithmic partition function consists of a dominant non-singular term $a(J)\beta L^{d}$ and a dimensionless singular term $\tilde{S}(\Delta gL^{1/\nu})$. The leading non-singular term obscures the critical behavior, making it difficult to extract the singular part directly from the partition function. More interestingly, as shown in Fig. 1, the curves of $S^{(2)}$ for different system sizes intersect at the critical point, indicates that the RTE is independent of system size, providing strong numerical evidence that the leading non-singular contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel exactly in the definition of RTE. This cancellation isolates the dimensionless singular term. Moreover, the DRTE displays a more pronounced divergence near the QCPs, further highlighting its sensitivity to critical fluctuations.
@@
According to the finite-size scaling hypothesis  [18, 17], any singular quantity near a critical point follows a scaling form as given in Eq.(1). In particular, the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
$S^{(n)}(\Delta g,L)=\tilde{S}(\Delta gL^{1/\nu})$ (5)
@@
$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
where $\tilde{S}^{\prime}(x)$ is $\partial\tilde{S}(x)/\partial x$ and $x$ is a general parameter.
Figure 1: (a,c) Second-order RTE and its derivative DRTE, and (b,d) corresponding data collapse analysis near the QCP of the bilayer Ising–Heisenberg model [(2+1)D Ising universality]. We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Results for the dimerized anisotropic Heisenberg model with $\Delta=0.9$ [(2+1)D O(2) universality], using $J_{2}=2.1035$ and tuning $J_{1}$. (i–m) Results for the dimerized isotropic Heisenberg model with $\Delta=1$ [(2+1)D O(3) universality], using $J_{2}=1$ and tuning $J_{1}$. To achieve good data collapse, we include a subleading correction: $Q(g,L)=L^{k}[\tilde{S}((g-g_{c})L^{1/\nu})+bL^{-\omega}]$. $\omega$ is close to 1, and to reduce the numerical instability caused by increasing the number of fitting parameters, we set $\omega=1$. For $Q=S^{(2)}$, $k=0$. $k=1/\nu$ when $Q$ is DRTE.
In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)D O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
Ising transition.— We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [43]. The Hamiltonian is given by:
@@
where $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ labels the two layers. $J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction. For $J/J_{z}\ll 1$, the ground state exhibits $\mathbb{Z}_{2}$ symmetry-breaking Ising order. In the opposite limit, $J/J_{z}\gg 1$, the system forms interlayer spin-singlet states. The QCP separating two phases is at $J/J_{z}=3.045(2)$ and belongs to the (2+1)D Ising universality class [43].
We set $J=3.045$ and use $J_{z}$ as the tuning parameter, studying the second-order RTE and DRTE near the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Figs. 1(a) and (c). The RTE curves for different system sizes intersect at the QCP, while the DRTE exhibits a clear peak at the critical point. We fit the scaling forms in Eqs. (5) and (6) to the RTE and DRTE data, respectively. The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.62(1)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE. These values are agreement with the known QCP $J_{z,c}=1$  [43] and the (2+1)D Ising universality class value $\nu=0.63012$  [44, 45, 43]. The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$, obtained from data collapse, are displayed in Figs. 1(b) and (d), respectively.
@@
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction with anisotropy parameter $\Delta$. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $0<\Delta<1$, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $0<\Delta<1$/ $\Delta>1$ / $\Delta=1$  [46, 47, 48].
@@
We evaluate the RTE and its derivative near the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality), with results shown in Figs. 1(e,g) and (i,k), respectively. In each case, $J_{2}$ is fixed and $J_{1}$ is tuned across the transition, as detailed in the caption. Similar to the Ising case, the RTE curves for different system sizes intersect at the QCPs, and their derivatives exhibit clear peaks at these points.
@@
Figure 2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$. The inset shows data on the right side of the transition point. (b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model. The raw values of the partition function derivative vary significantly across system sizes (see SM), making direct comparison difficult. To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
We apply the scaling forms in Eqs. (5) and (6) to data near the (2+1)D O(2) and O(3) QCPs. For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.999(1)$, $\nu=0.66(2)$. These values are consistent with the known critical parameters $J_{1,c}=1$ and $\nu=0.6703$ [44, 48].
For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.70(1)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.71(2)$. These values are consistent with the known critical parameters $J_{1,c}=0.52337$ and $\nu=0.7073$ [44, 46, 47]. The universal scaling functions near the O(2) and O(3) QCPs, obtained from data collapse, are shown in Figs. 1(f,h) and (j,m), respectively.
@@
First-order and weakly first-order transitions.— Having established the universal scaling behavior at (2+1)D QCPs, we now turn to first-order phase transitions, with a focus on weakly first-order transitions, which presents significant challenges for traditional local physical quantities.
@@
We first study the 2D checkerboard $J$-$Q$ model ( see SM for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [49]. The DRTE around this transition is plotted in Fig. 2 (a). A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side as $L\to\infty$. This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above). For comparison, we also calculate the derivative of logarithmic partition functions, as shown in Figs. 2 (b) and (c). These curves appear much smoother across the transition compared with DRTE, due to the dominance of a non-singular, volume-law term. The phase transition of this model is obviously first-order according to the previous works [49, 29], thus we see that the curves in (b) and (c) have very small kinks around the phase transition point, especially for larger sizes.
@@
The phase transition between AF Néel order and valence-bond solid (VBS) phase in 2D quantum magnets can appear (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [4, 5]. $J-Q_{2}$ and $J-Q_{3}$ models are prototypical examples that realize such transitions in 2D quantum spin systems [50, 51]. These models are free of the sign problem, making them ideal for QMC simulations. Extensive QMC studies have shown that the phase transitions of the JQJ-Q models on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [50, 51, 52, 53, 54, 55]. However, the critical exponents observed in these models do not satisfy the conformal bootstrap bounds expected for SO(5) critical point [56]. Recent works have increasingly suggested that this DQC scenario may instead correspond to a weakly first-order transition or as a nonunitary CFT  [7, 11, 57, 58, 12, 59, 60, 61, 8, 15, 62, 63, 64, 65, 16]. In the following, we show that the DRTE provides a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see SM for models’ details) are indeed first-order, even in rather small systems (up to $L=56$).